# ScoutGPT: Capturing Player Impact from Team Action Sequences Using GPT-Based Framework

### v2

@@ Abstract.
Transfers play a pivotal role in shaping a football club’s success, yet forecasting whether a transfer will succeed remains difficult due to the strong context-dependence of on-field performance. Existing evaluation practices often rely on static summary statistics or post-hoc value models, which fail to capture how a player’s contribution adapts to a new tactical environment or different teammates. To address this gap, we introduce ScoutGPT, a player-conditioned, value-aware next-event prediction model built on a GPT-style autoregressive transformer. Our model treats match play as a sequence of discrete tokens, jointly learning to predict the next on-ball action’s type, location, timing, and its estimated residual On-Ball Value (rOBV) based on the preceding context and player identity. A key contribution of this framework is the ability to perform counterfactual simulations. By substituting learned player embeddings into new event sequences, we can simulate how a player’s behavioral distribution and value profile would change when placed in a different team or tactical structure. Evaluated on five seasons of Premier League event data, ScoutGPT outperforms existing sequence-based baselines in next-event prediction accuracy and spatial precision. Furthermore, we demonstrate the model’s practical utility for transfer analysis through case studies—such as comparing striker performance across different systems and identifying stylistic replacements for specific roles—showing that our approach provides a principled method for evaluating transfer fit.
@@ 1. Introduction
Transfers play a pivotal role in shaping a football club’s sporting and financial future, often determining its long-term success or failure. The importance of the market keeps rising: according to FIFA, the 2025 mid-year window alone saw a record USD 9.76 billion in transfer fees, with England the top-spending association  (FIFA2025Transfers). Nevertheless, forecasting whether a transfer will succeed remains difficult. Post-transfer performance is influenced by multiple interacting factors, including the competitive level of the destination league, the tactical philosophy of the new coaching staff, the player’s assigned role within team structure, squad hierarchy and chemistry, physical adaptation and workload patterns, and even cultural and environmental acclimatization (Berlinschi2013; Jarjabka2024). These elements interact in ways that are nonlinear and often opaque, making simple evaluation based on past performance insufficient.
@@
A substantial line of research has focused on evaluating how much an on-ball action contributes to a team’s chance of scoring or preventing goals. Decroos et al. (DecroosVAEP2021) introduced VAEP, a framework that estimates the change in scoring and conceding probabilities induced by each action, demonstrating that value is not inherent to the action itself but emerges from its context within the game state. Similarly, Expected Threat (xT) introduced a spatial scoring-potential surface, emphasizing how ball progression into certain pitch zones increases goal probability  (Singh2019xT). On-Ball Value (OBV) assigns value to each on-ball action based on how it changes the expected outcome of the current possession sequence  (StatsBomb2021OBV). More recently, Expected Possession Value (EPV) extends these ideas by estimating the value of entire possessions and modeling how present states imply future scoring chances (Fernandez2019EPV). These methods have been influential in highlighting that action impact is fundamentally contextual. However, value estimation is typically applied as a post-hoc layer on completed event sequences; the value of an action is evaluated after it happens, rather than co-learned with the sequential process that generates actions. As a result, these models capture what was valuable, but not how that value would change when a player operates in a different tactical environment.
@@
Another line of research has focused on learning player representations that capture style, role, and behavioral tendencies. Early approaches constructed player profiles from aggregated event statistics or possession chain summaries, while later work introduced embedding-based methods that learn latent representations from match event sequences or network structures (Pappalardo2019; DecroosVector2019). Such embeddings support tasks such as similarity search, role classification, and recruitment shortlist generation, and they have proven effective for identifying players with comparable tactical functions. More recent work incorporates relational structure or interaction patterns, for example, by modeling passing networks or co-occurring on-ball behaviors, in order to represent how a player’s contributions depend on teammates and system shape  (YilmazEmbedding2022; Rovshitz2024TransformerSoccer; adjileye2024risingballer). However, these representations are generally static: they summarize how a player has behaved in their previous context, but they do not model how that behavior would adapt when the surrounding tactical environment changes. Consequently, while embedding approaches help answer the question of who a player is, they do not address the core transfer-fit question of what this player would do in a different team, under different roles and constraints.
@@
Recent work has approached match play from a sequential modeling perspective, treating event streams as structured sequences rather than isolated observations. Early next-event prediction models, such as Seq2Event (Simpson2022Seq2event), demonstrated that the unfolding flow of possession can be captured autoregressively, learning to predict the next action based on preceding context and thereby revealing regularities in tactical organization. Extending this idea, neural marked spatio-temporal point process models (NMSTPP)  (Yeung2025NMSTPP) jointly model the type, location, and timing of future actions, emphasizing that match dynamics are fundamentally spatio-temporal rather than purely categorical. More recently, Large Event Models (LEM)  (MendesNeves2026LEM) have advanced this autoregressive view toward a generative framing in which full match progressions can be simulated from arbitrary game states, moving toward foundation-model-style applications such as scenario exploration and tactical evaluation. However, existing sequence-based and large-event models typically do not explicitly condition on player identity and generally do not jointly learn action value within the generative process. As a result, they describe how play unfolds in aggregate, but do not account for how the presence of a specific player would alter the trajectory and value of actions within a new tactical environment—precisely the core problem in evaluating transfer fit.
@@
Figure 1. Overview of the ScoutGPT framework. Our nanoGPT-based Transformer model autoregressively predicts event tokens, enabling counterfactual ‘what-if’ simulations. For instance, replacing Darwin Núñez with Alexander Isak could alter actions (e.g., pass/shot) or modify the same action with a different location, outcome, or on-ball value (OBV). A video example of this sequence can be seen at the 23-second mark in the linked video clip.
@@
To address this gap, we introduce ScoutGPT, a player-conditioned, value-aware next-event prediction model built on a GPT-style autoregressive transformer. Figure 1 provides an overview of the ScoutGPT framework, illustrating how our model autoregressively predicts event tokens, enabling counterfactual ’what-if’ simulations. The model jointly predicts the next on-ball action and its estimated on-ball value, with player identity provided as an input through learned embeddings that influence the prediction of all downstream event attributes, including action type, spatial location, timing, outcome, and action-level value. Crucially, the learned player embeddings can be substituted into new team event sequences, enabling counterfactual simulation of how a player’s action distribution and value profile would change when placed in a different tactical environment. This provides a principled framework for context-aware transfer fit evaluation, which existing action-value models, player representations, and sequence-based event models do not directly support.
@@ 2.3. Model Architecture
follows a decoder-only Transformer architecture directly adapted from the NanoGPT implementation of the GPT model by Karpathy  (Karpathy2022nanoGPT). The network consists of a stack of causal self-attention decoder blocks, each comprising multi-head attention and position-wise feedforward layers with residual connections and layer normalization. Token embeddings are applied through a single shared embedding matrix over the unified vocabulary described in Section 2.1, and learned positional embeddings are added to preserve event order within each episode. The final layer ties the output projection weights to the input embedding matrix, enabling efficient next-token prediction across all event attributes.
@@ 3.4. Main Results
Notably, our model improves event type prediction accuracy compared to both baselines (82.91% vs. 80.42% for the LEM Transformer and 74.07% for LEM), indicating that conditioning on player identity and episode context leads to more accurate modeling of action selection. In addition, our model substantially reduces spatial prediction error, demonstrating improved understanding of how play unfolds on the field. While the LEM Transformer obtains the lowest rOBV error (0.008), ScoutGPT achieves a competitive second-best result (0.009).

### v3

# EventGPT: Capturing Player Impact from Team Action Sequences Using GPT-Based Framework

@@ Abstract.
Transfers play a pivotal role in shaping a football club’s success, yet forecasting whether a transfer will succeed remains difficult due to the strong context-dependence of on-field performance. Existing evaluation practices often rely on static summary statistics or post-hoc value models, which fail to capture how a player’s contribution adapts to a new tactical environment or different teammates. To address this gap, we introduce EventGPT, a player-conditioned, value-aware next-event prediction model built on a GPT-style autoregressive transformer. Our model treats match play as a sequence of discrete tokens, jointly learning to predict the next on-ball action’s type, location, timing, and its estimated residual On-Ball Value (rOBV) based on the preceding context and player identity. A key contribution of this framework is the ability to perform counterfactual simulations. By substituting learned player embeddings into new event sequences, we can simulate how a player’s behavioral distribution and value profile would change when placed in a different team or tactical structure. Evaluated on five seasons of Premier League event data, EventGPT outperforms existing sequence-based baselines in next-event prediction accuracy and spatial precision. Furthermore, we demonstrate the model’s practical utility for transfer analysis through case studies—such as comparing striker performance across different systems and identifying stylistic replacements for specific roles—showing that our approach provides a principled method for evaluating transfer fit.
@@ 1. Introduction
Transfers play a pivotal role in shaping a football club’s sporting and financial future, often determining its long-term success or failure. The importance of the market keeps rising: according to FIFA, the 2025 mid-year window alone saw a record USD 9.76 billion in transfer fees, with England the top-spending association  (FIFA, 2025). Nevertheless, forecasting whether a transfer will succeed remains difficult. Post-transfer performance is influenced by multiple interacting factors, including the competitive level of the destination league, the tactical philosophy of the new coaching staff, the player’s assigned role within team structure, squad hierarchy and chemistry, physical adaptation and workload patterns, and even cultural and environmental acclimatization (Berlinschi et al., 2013; Jarjabka et al., 2024). These elements interact in ways that are nonlinear and often opaque, making simple evaluation based on past performance insufficient.
@@
A substantial line of research has focused on evaluating how much an on-ball action contributes to a team’s chance of scoring or preventing goals. Decroos et al. (Decroos et al., 2020) introduced VAEP, a framework that estimates the change in scoring and conceding probabilities induced by each action, demonstrating that value is not inherent to the action itself but emerges from its context within the game state. Similarly, Expected Threat (xT) introduced a spatial scoring-potential surface, emphasizing how ball progression into certain pitch zones increases goal probability  (Singh, 2019). On-Ball Value (OBV) assigns value to each on-ball action based on how it changes the expected outcome of the current possession sequence  (StatsBomb, 2021). More recently, Expected Possession Value (EPV) extends these ideas by estimating the value of entire possessions and modeling how present states imply future scoring chances (Fernández et al., 2019). These methods have been influential in highlighting that action impact is fundamentally contextual. However, value estimation is typically applied as a post-hoc layer on completed event sequences; the value of an action is evaluated after it happens, rather than co-learned with the sequential process that generates actions. As a result, these models capture what was valuable, but not how that value would change when a player operates in a different tactical environment.
@@
Another line of research has focused on learning player representations that capture style, role, and behavioral tendencies. Early approaches constructed player profiles from aggregated event statistics or possession chain summaries, while later work introduced embedding-based methods that learn latent representations from match event sequences or network structures (Pappalardo et al., 2019; Decroos and Davis, 2020). Such embeddings support tasks such as similarity search, role classification, and recruitment shortlist generation, and they have proven effective for identifying players with comparable tactical functions. More recent work incorporates relational structure or interaction patterns, for example, by modeling passing networks or co-occurring on-ball behaviors, in order to represent how a player’s contributions depend on teammates and system shape  (Yılmaz and Öğüdücü, 2022; Rovshitz and Puzis, 2024; Adjileye, 2024). However, these representations are generally static: they summarize how a player has behaved in their previous context, but they do not model how that behavior would adapt when the surrounding tactical environment changes. Consequently, while embedding approaches help answer the question of who a player is, they do not address the core transfer-fit question of what this player would do in a different team, under different roles and constraints.
@@
Recent work has approached match play from a sequential modeling perspective, treating event streams as structured sequences rather than isolated observations. Early next-event prediction models, such as Seq2Event (Simpson et al., 2022), demonstrated that the unfolding flow of possession can be captured autoregressively, learning to predict the next action based on preceding context and thereby revealing regularities in tactical organization. Extending this idea, neural marked spatio-temporal point process models (NMSTPP)  (Yeung et al., 2025) jointly model the type, location, and timing of future actions, emphasizing that match dynamics are fundamentally spatio-temporal rather than purely categorical. More recently, Large Event Models (LEM)  (Mendes-Neves et al., 2026) have advanced this autoregressive view toward a generative framing in which full match progressions can be simulated from arbitrary game states, moving toward foundation-model-style applications such as scenario exploration and tactical evaluation. However, existing sequence-based and large-event models typically do not explicitly condition on player identity and generally do not jointly learn action value within the generative process. As a result, they describe how play unfolds in aggregate, but do not account for how the presence of a specific player would alter the trajectory and value of actions within a new tactical environment—precisely the core problem in evaluating transfer fit.
@@
Figure 1. Overview of the EventGPT framework. Our nanoGPT-based Transformer model autoregressively predicts event tokens, enabling counterfactual ‘what-if’ simulations. For instance, replacing Darwin Núñez with Alexander Isak could alter actions (e.g., pass/shot) or modify the same action with a different location, outcome, or on-ball value (OBV). A video example of this sequence can be seen at the 23-second mark in the linked video clip.
@@
To address this gap, we introduce EventGPT, a player-conditioned, value-aware next-event prediction model built on a GPT-style autoregressive transformer. Figure 1 provides an overview of the EventGPT framework, illustrating how our model autoregressively predicts event tokens, enabling counterfactual ’what-if’ simulations. The model jointly predicts the next on-ball action and its estimated on-ball value, with player identity provided as an input through learned embeddings that influence the prediction of all downstream event attributes, including action type, spatial location, timing, outcome, and action-level value. Crucially, the learned player embeddings can be substituted into new team event sequences, enabling counterfactual simulation of how a player’s action distribution and value profile would change when placed in a different tactical environment. This provides a principled framework for context-aware transfer fit evaluation, which existing action-value models, player representations, and sequence-based event models do not directly support.
@@ 2.3. Model Architecture
follows a decoder-only Transformer architecture directly adapted from the NanoGPT implementation of the GPT model by Karpathy  (Karpathy, 2022). The network consists of a stack of causal self-attention decoder blocks, each comprising multi-head attention and position-wise feedforward layers with residual connections and layer normalization. Token embeddings are applied through a single shared embedding matrix over the unified vocabulary described in Section 2.1, and learned positional embeddings are added to preserve event order within each episode. The final layer ties the output projection weights to the input embedding matrix, enabling efficient next-token prediction across all event attributes.
@@ 3.4. Main Results
Notably, our model improves event type prediction accuracy compared to both baselines (82.91% vs. 80.42% for the LEM Transformer and 74.07% for LEM), indicating that conditioning on player identity and episode context leads to more accurate modeling of action selection. In addition, our model substantially reduces spatial prediction error, demonstrating improved understanding of how play unfolds on the field. While the LEM Transformer obtains the lowest rOBV error (0.008), EventGPT achieves a competitive second-best result (0.009).